# Detecting AI Flaws: Target-Driven Attacks on Internal Faults in Language Models

### v1

@@ 5.2 Adversarial Attacks on LLMs
Despite the rapid adoption of applications built on aligned large language models (LLMs), users have discovered that carefully phrased prompts can still elicit malicious responses from these models. Consequently, addressing vulnerabilities in LLMs has become essential for enhancing their robustness and safety. There are three primary methods of attack. The first involves manually crafting phishing queries. A notable example is DAN (walkerspider 2022), which uses hand-designed prompts to exploit vulnerabilities in online chatbots powered by aligned LLMs. The second method is optimization-based. GCG (Zou et al. 2023) automates the prompt generation process by utilizing gradient information from open-source LLMs to guide the search for optimal tokens, potentially leading to unexpected responses. PAL (Sitawarin et al. 2024) proposes transferring knowledge from open-source LLMs to closed-source models such as GPT-3.5, enabling attacks on black-box models without needing access to their gradients. The third method accelerates prompt generation without using the target models’ gradients. Advprompter (Paulus et al. 2024) employs a different LLM as the prompt generator to create adversarial suffixes based on the original prompts. BEAST (Sadasivan et al. 2024) further optimizes the balance between attack speed and success rate, enabling the attack to be performed more quickly on a single GPU.
@@ 6 Conclusion
In this paper, we propose a target-driven attacking paradigm that employs a novel approach involving another LLM, named ToxDet, to generate potential questions and preliminary answers designed to elicit inappropriate responses from the target model. We optimize ToxDet using reinforcement learning, leveraging the target LLM itself as the reward model, thereby eliminating the need for the inclusion of a separate reward model. Our primary focus is on interacting with gray-box target LLMs by utilizing probability information; however, the results can also be effectively transferred to black-box models like GPT-4o, achieving favorable outcomes. Extensive experimental results on the AdvBench and HH-Harmless datasets demonstrate the effectiveness of our proposed methods, revealing vulnerabilities in LLMs and providing a valuable resource for researchers aiming to enhance model safety.
Ethical Statement
@@
This paper introduces an automated approach for generating questions and preliminary answers designed to manipulate LLMs into producing outputs that diverge from human preferences, intentions, or values. We posit that, similar to previous attacking research, this work is unlikely to cause immediate harm. Instead, it aims to foster the development of more effective defensive strategies, ultimately leading to more robust, safe, and aligned LLMs. We hope that by highlighting the vulnerabilities of LLMs, this research will attract attention from both academia and industry, prompting the creation of stronger defenses and more rigorous safety measures, thereby enabling LLMs to better serve society.
Appendix
Appendix A Necessity of Preliminary Answers
@@
A primary concern is why we add a preliminary answer instead of a potential question directly. We found that using the question itself inherently optimizes the prompt in a way that has minimal impact on the target model. Therefore, we propose the use of a preliminary answer. To intuitively demonstrate the necessity of employing preliminary answers, we present comparative examples. Specifically, we provide responses from ToxDet with and without preliminary answers, as well as the corresponding responses from the target LLM, in Fig. 3. In this table, we highlight the toxic components of ToxDet’s responses in red and those of the target LLM in blue. The comparison clearly shows that responses from the target LLM, when preliminary answers are used, exhibit more toxic elements, whereas responses from ToxDet contain more toxic and misleading parts. Therefore, we conclude that incorporating a preliminary answer is beneficial.
@@
Figure 3: Examples of ToxDet responses with and without preliminary answers, along with the corresponding target LLM outputs. We use the Mistral-7b model without fine-tuning as ToxDet and Llama3-8b as the target LLM. Toxic content is indicated by color: red for ToxDet responses and blue for target LLM outputs.
Appendix B Perplexity Evalutaion
@@
We report the perplexity (PPL) of the generated content conditioned on the prompts. For AdvPrompter and other comparison methods, PPL is measured based on the generated suffix by the attack conditioned on the prompts. In contrast, for our method, PPL is measured based on the generated precedent content, i.e., the prompt and the preliminary answer conditioned on the target response.
@@
Theoretically, these two measures are not directly comparable. However, in practice, the generated precedent content generally has a longer sequence length than the suffix generated by AdvPrompter and other models. This difference in sequence length could result in a higher PPL for the precedent content of our method compared to the PPL of the suffix in other methods under the same conditions. Therefore, this comparison remains fair.
@@ Appendix C System Prompts
The system prompt for querying ToxDet is as follows:
@@ Appendix D Failure Cases
Then, we provide some failure cases, as shown in Fig. 5, where using the generated precedent content from ToxDet fails to attack the target LLM. There are two types of failure cases. The first type is evident, where the target LLM responds with keywords like “I cannot …,” which can be easily detected by keyword evaluation, as highlighted in orange. The second type of failure case occurs when the target LLM does not provide the desired toxic response. Instead, it responds with content that suggests how to act legally, as highlighted in green. This type of failure case can only be detected by GPT-4 evaluation, which requires prompting GPT-4 to understand the response content.
Figure 5: Cases of generated precedent content from ToxDet fail to attack the target LLM. We use the fine-tuned Mistral-7b for ToxDet and Llama3 for the target LLM as examples. Orange cases indicate instances that can be easily detected by keyword lists, while green cases indicate instances that can only be detected by GPT-4 by understanding the contents.
@@ Appendix E More Case Study
In this section, we provide more examples of attacks on both gray-box and black-box target LLMs, as illustrated in Fig. 6 and Fig. 7.
Figure 6: Cases of successfully attacking the gray-box target LLMs. Original target responses starting with “Sure” are sourced from AdvBench, while the others are from HH-Harmless.
Figure 7: Cases of successfully transfer attacking the black-box target LLMs with ToxDet interactively trained with Llama3-8b.

### v2

@@ 5.2 Adversarial Attacks on LLMs
Despite the rapid adoption of applications built on aligned large language models (LLMs), users have discovered that carefully phrased prompts can still elicit malicious responses from these models. Consequently, addressing vulnerabilities in LLMs has become essential for enhancing their robustness and safety. There are three primary methods of attack. The first involves manually crafting phishing queries. A notable example is DAN (walkerspider, 2022), which uses hand-designed prompts to exploit vulnerabilities in online chatbots powered by aligned LLMs. The second method is optimization-based. GCG (Zou et al., 2023) automates the prompt generation process by utilizing gradient information from open-source LLMs to guide the search for optimal tokens, potentially leading to unexpected responses. PAL (Sitawarin et al., 2024) proposes transferring knowledge from open-source LLMs to closed-source models such as GPT-3.5, enabling attacks on black-box models without needing access to their gradients. The third method accelerates prompt generation without using the target models’ gradients. AdvPrompter (Paulus et al., 2024) employs a different LLM as the prompt generator to create adversarial suffixes based on the original prompts. BEAST (Sadasivan et al., 2024) further optimizes the balance between attack speed and success rate, enabling the attack to be performed more quickly on a single GPU.
@@ 6 Conclusion
We introduce the red-teaming Attacker with Target Toxic Answers (Atoxia), a language model designed to generate adversarial questions and answer openings to induce an under-testing LLM to output inappropriate or harmful responses. Atoxia is optimized using reinforcement learning, where the under-testing LLM itself serves as the reward model, eliminating the need for an independent RM training process, which allows Atoxia to effectively detect safety vulnerabilities of the under-testing LLM. While our primary effort is on gray-box under-testing LLMs, where probability information is accessible, our approach also generalizes well to black-box models like GPT-4o, achieving similarly robust results. Comprehensive experimental evaluations on the AdvBench and HH-Harmless datasets validate the effectiveness of our method, which successfully uncovers vulnerabilities in multiple LLMs. These discoveries offer valuable insights for researchers working on enhancing the safety and robustness of LLMs, providing a proactive method for identifying and mitigating risks before they can be exploited in real-world scenarios.
Acknowledgments
This work was supported by the Guangxi Key R&D Project (No. AB24010167), the Project (No. 20232ABC03A25), Futian Healthcare Research Project (No.FTWS002).
Limitations
While our proposed Atoxia framework demonstrates promising results in detecting toxic content across both open-source and black-box models, it has several limitations. One key challenge is the computational intensity of the reinforcement learning optimization process, which becomes particularly demanding when applied to larger models. Due to limited computational resources, we are currently unable to scale our approach effectively to larger models. Additionally, given the rapid advancements in LLMs, our framework has been tested on the currently popular models, such as Llama3 and GPT-4o. However, as newer models emerge, we cannot guarantee the same level of performance, as our detection framework may need further adaptation to handle the evolving architectures and behaviors of future LLMs. This highlights the need for continuous evaluation and updating of our approach.
Ethical Considerations
@@ Acknowledgments
This paper introduces an automated approach for generating questions and answer openings designed to manipulate LLMs into producing outputs that diverge from human preferences, intentions, or values. We posit that, similar to previous attacking research, this work is unlikely to cause immediate harm. Instead, it aims to foster the development of more effective defensive strategies, ultimately leading to more robust, safe, and aligned LLMs. We hope that by highlighting the vulnerabilities of LLMs, this research will attract attention from both academia and industry, prompting the creation of stronger defenses and more rigorous safety measures, thereby enabling LLMs to better serve society.
Appendix A Necessity of Answer Opening
@@
A primary concern is why we add an answer opening ${\bm{y}}_{\text{pre}}$ instead of a potential question directly. We found that using the question itself inherently optimizes the prompt in a way that has minimal impact on the target model. Therefore, we propose the use of an answer opening. To intuitively demonstrate the necessity of employing answer openings, we present comparative examples. Specifically, we provide responses from Atoxia with and without answer openings, as well as the corresponding responses from the under-testing (UT) LLM, in Fig. 3. In this table, we highlight the toxic components of Atoxia’s responses in red and those of the under-testing LLM in blue. The comparison clearly shows that responses from the under-testing LLM, when answer openings are used, exhibit more toxic elements, whereas responses from Atoxia contain more toxic and misleading parts. Therefore, we conclude that incorporating an answer opening is beneficial.
@@
Figure 3: Examples of Atoxia responses with and without answer openings, along with the corresponding under-testing (UT) LLM outputs. We use the Mistral-7b model without fine-tuning as Atoxia and Llama3-8b as the under-testing LLM. Toxic content is indicated by color: red for Atoxia responses and blue for under-testing LLM outputs.
Appendix B Perplexity Evaluation
@@
We report the perplexity (PPL) of the generated content conditioned on the prompts. For AdvPrompter and other comparison methods, PPL is measured based on the generated suffix by the attack conditioned on the prompts. In contrast, for our method, PPL is measured based on the generated adversarial question and answer opening conditioned on the target answer.
@@
Theoretically, these two measures are not directly comparable. However, in practice, the generated question and answer opening generally have a longer sequence length than the suffix generated by AdvPrompter and other models. This difference in sequence length could result in a higher PPL for the generated content of our method compared to the PPL of the suffix in other methods under the same conditions. Therefore, this comparison remains fair.
@@ Appendix C System Prompts
The system prompt for querying Atoxia is as follows:
@@ Appendix D Failure Cases
Then, we provide some failure cases, as shown in Fig. 5, where using the generated content from Atoxia fails to attack the under-testing LLM. There are two types of failure cases. The first type is evident, where the under-testing LLM responds with keywords like “I cannot …,” which can be easily detected by keyword evaluation, as highlighted in orange. The second type of failure case occurs when the under-testing LLM does not provide the desired toxic response. Instead, it responds with content that suggests how to act legally, as highlighted in green. This type of failure case can only be detected by GPT-4 evaluation, which requires prompting GPT-4 to understand the response content.
Figure 5: Cases of generated content from Atoxia fail to attack the under-testing LLM. We use the fine-tuned Mistral-7b for Atoxia and Llama3 for the under-testing LLM as examples. Orange cases indicate instances that can be easily detected by keyword lists, while green cases indicate instances that can only be detected by GPT-4 by understanding the contents.
@@ Appendix E More Case Study
In this section, we provide more examples of attacks on both gray-box and black-box under-testing LLMs, as illustrated in Fig. 6 and Fig. 7.
Figure 6: Cases of successfully attacking the gray-box under-testing LLMs. Original target answers starting with “Sure” are sourced from AdvBench, while the others are from HH-Harmless.
Figure 7: Cases of successfully transfer attacking the black-box under-testing LLMs with Atoxia interactively trained with Llama3-8b.